# Bridge indices of spatial graphs and diagram colorings

### v1

@@ 4.2 Meridional rank
The fundamental group of the complement of a spatial graph $G$ in $S^{3}$ can be computed via the Wirtinger algorithm in a similar way to the computations for knots and links. Each arc of a spatial graph diagram gives a generator, and each crossing gives a relation of the form $x_{i}x_{j}x^{-1}_{i}x^{-1}_{k}$. At a vertex, we have a new type of relation of the form $x_{i_{1}}^{\epsilon_{1}}x_{i_{2}}^{\epsilon_{2}}\cdots x_{i_{n}}^{\epsilon_{n}}$, where the generators are listed in order as we go clockwise around a vertex, and $\epsilon_{i}$ is 1 (resp. $-1$) if the arc is directed into the vertex (resp. directed out of the vertex). See Figure 9.
@@
The technique of bounding the number of handles of knotted objects by the meridional rank of the fundamental group has been used before in other settings such as classical links [baader2021coxeter] and surface-links [joseph2024meridional]. We now adapt the method to the setting of spatial graphs.
In [livingston1995knotted] Livingston defined the vertex constant group $\pi^{*}$ to be the quotient of $\pi_{1}(S^{3}\backslash G)$ defined by setting all meridians at each vertex equal to each other. We are interested in bounding the rank of the vertex constant group, where our generating set consists of meridians. We achieve this by searching for a consistent labeling of our diagram by reflections in a Coxeter group $C(\Delta)$. This will produce a surjective homomorphism from $\pi^{*}$ to $C(\Delta)$.
@@
Recall that a presentation of a Coxeter group $C(\Delta)$ can be obtained from a weighted graph $\Delta$ as follows. Label the vertices of $\Delta$ as $x_{1},x_{2},\cdots,x_{n}$, which will be the generators. If there is an edge with weight $w$ connecting $x_{i}$ to $x_{j}$, then we have a relation $(x_{i}x_{j})^{w}$ in the presentation. Additionally, the relations also include $x_{i}^{2},$ for $i=1,2,\cdots,n.$
@@
Lemma 4.6 ([felikson2010reflection, Lem. 2.1]).
If all edge weights of $\Delta$ are at least two, then the reflection rank equals the minimal number of reflections needed to generate $C(\Delta).$
@@
A labeling of arcs in a spatial graph diagram $D$ is consistent if at each crossing where the overstrand is labeled with an element $g$ and the two understrands are labeled $h$ and $k$, we have that $ghg=k$. We say that the label generates the group if every element of $C(\Delta)$ can be written as a product of the elements that appear as labels along with their inverses.
@@
If a diagram $D$ representing a spatial graph $G$ can be labeled by $n$ reflections from a Coxeter group $C(\Delta)$ with reflection rank $n$, then $\beta(G)\geq n.$
@@
In the same spirit as Proposition 3.4 and Proposition 4.4, the labels on the upper pods and maxima determine all the other labels. The difference is that we use a labeling by Coxeter groups in this setting. Thus, these labels generate $C(\Delta)$. We then appeal to Lemma 4.6 to conclude that we need at least $n$ reflections to generate $C(\Delta)$. ∎
@@ 5 Computations
In this section, we use our techniques to estimate or provide exact values for the bridge indices of many different examples of knotted spatial graphs. In particular, although the following statement also follows from other techniques, our coloring and meridional rank techniques give short arguments, as we will see in Corollary 5.9, Corollary 5.13, and Corollary 5.18.
See 1.2
Refer to Section 2.1 for a review of the spatial graph terminology we use throughout this section. We break into various subsections according to the type of construction, but start first with a relatively simple example.
@@
As a warmup, Figure 10 gives an almost unknotted handcuff graph (left) and an almost unknotted $\Theta$-graph (right). Both of these well-known examples – for instance the $\Theta$-graph is the Kinoshita graph [kinoshita1958alexander, kinoshita1972elementary] – were constructed by adding an edge to an unlink, and both have bridge index $2$. These diagrams are already in bridge position demonstrating this index, but this can also be verified by finding $2$-Wirtinger colorings of the diagrams. The lower bound comes from the fact that there is a nontrivial coloring by the Alexander quandle of order four from Section 4.1. Note that the number of crossings in each column of twists can be systematically increased and we still retain the consistency of the coloring. By systematic, we mean that if there are three crossings originally, we can keep adding a multiple of three crossings.
@@
Figure 10: A quandle colored almost unknotted handcuff graph (left) and a quandle colored almost unknotted $\Theta$-graph (right), both with bridge index $2$.
@@ 5.1 Python code
We implemented our algorithm in Python to compute upper bounds for the Wirtinger numbers of arbitrary $\Theta_{n}$-graphs ($n\leq 26)$. Our algorithm generalizes coloring algorithms used to compute bounds for the Wirtinger numbers for knots and links. We refer readers to [blair2020wirtinger, lee2024widths] for details on the knot and link cases. Our Python code can be found at https://github.com/hanhv/graph-wirt. While the full detail can be found in the URL, we also give a summary here.
@@
A Gauss code for an $m$-component link is a list of $m$ lists of positive and negative integers such that for all $k\neq 0,|k|\leq m$, both $k$ and $-k$ appear once each among the lists.
@@
Figure 11: Example of a Gauss code. This diagram is 3-Wirtinger colorable using the seed [{a1, b1, c1}, {a2, b2, c2}, {-8, 9, -7}], which is colored pink.
@@
First we extract two-component link diagrams from the link table provided by SnapPy [SnapPy]. Each of these links takes the form of a Gauss code made up of two lists of numbers. Selecting a two-component link, by “singularizing” we turn two crossings between components into 4-valent vertices. Here, a pair of crossings $(x,y)$ is considered singularizable if both $x$ and $y$ appear in one component, while $-x$ and $-y$ appear in the other component. We only choose singularizable pairs that are of distance at least four apart (this is optional), ensuring that each edge of the graph contains at least four crossings. By singularizing a pair of crossings $i$ and $j$, we split each component of the link, viewed as a cyclic list, into two parts with endpoints at $i$ and $j$. This has the effect of turning the two-component link into a $\Theta_{4}$-graph.
@@
The program begins by identifying all strands in the Gauss code, which can be classified as pods (arcs in the neighborhood of a vertex), arcs connecting the endpoints of pods to undercrossings, or arcs connecting two undercrossings. It first sets $k=1$ and generates a list of seeds, which are combinations of $k$ strands, ensuring that at least one pod is included in each combination. (Our algorithm does not require this, but adding in this requirement will still produce a viable upper bound.) For each seed, it colors the arcs in the seed and creates a truncated version of the graph by removing any pods that are not part of the selected seed. The program then attempts to maximally extend this truncated graph through two types of coloring moves: one that extends from the endpoint of a pod and another that extends from an undercrossing if its corresponding overstrand is colored. If the entire truncated graph can be colored, the function returns $k$. If not, it tries the next seed. Should no combination successfully color the truncated graph, the algorithm increments $k$ and repeats the process.
@@
We extracted and computed the Wirtinger numbers for over 260,000 $\Theta_{4}$-graph diagrams.
@@ 5.2 Hara’s graphs and modifications
The family of almost unknotted $\Theta_{4}$-graphs given by Hara [hara1991symmetry, Fig. 2] all have bridge index $2$, which can be quickly verified by finding $2$-Wirtinger colorings of the diagrams. This is perhaps surprising, as these particular diagrams have many more local maxima and minima.
@@
For the sake of simplicity, we depict both families as coming from $\Theta$-graphs rather than $\Theta_{4}$-graphs, but both could be further extended by adding more edges in analogous ways. This would likely not change the bridge index. Note that with this change, we can (perhaps more appropriately) think of these families as modifications of a Kinoshita-Wolcott graph [kinoshita1958alexander, kinoshita1972elementary, wolcott1987knotting] (see [jang2016new, Fig. 1]).
@@
Let $G_{n}$ be the $\Theta$-graph obtained from Hara’s graphs by replacing the braids on two strands with braids on $n$ strands. Then the bridge index of $G_{n}$ is less than or equal to $n$.
@@
Let $G_{n}$ be the $\Theta_{3n}$-graph obtained from Hara’s graphs by taking $n$ parallel copies of each edge. Then the bridge index of $G_{n}$ is less than or equal to $n+1$.
@@ 5.3 Families from vertex sums
Take $n$ copies of the almost unknotted Suzuki $\Theta_{4}$-graph [suzuki1984almost], as shown in Figure 13. Consider the quandle ($(\{0,1,2\},\triangleright)$, where $x\triangleright y=2y-x\pmod{3}$. It can be shown that this quandle is homogeneous (see [furuki2024homogeneous], for example). In Figure 13 we give three consistent colorings of the graph by this quandle; then by permuting the colors, we have a total of nine possible colorings.
Taking the vertex sum of $n$ copies of the graph, we obtain a diagram with bridge index $n+1$. Using Proposition 4.4, we get that the number of colorings of the resulting vertex sum is $3^{n+1}$. Then by Proposition 4.3, the bridge index is precisely $n+1$. Thus we have shown the following.
@@
Let $G_{n}$ be the $\Theta_{4}$-graph obtained by taking the vertex sum of $n$ copies of Suzuki’s graph. Then the bridge index of $G_{n}$ is $n+1$.
Figure 13: For a Suzuki $\Theta_{4}$-graph, we see that the number of colorings where the strands next to a particular vertex are labeled by a fixed element of the dihedral quandle of order three is three. (Here we use colors to represent the labelings by the quandle.) Thus by permuting the colors, we have a total of nine colorings.
@@
The clasp move [simon1990minimally] (see [jang2016new, Fig. 3]), is a well-known move which preserves almost unknottedness of $\Theta$-graphs. We will use this to build a family of graphs as follows. First construct an almost unknotted $\Theta$-graph by performing the clasp move on a slightly isotoped trivial $\Theta$-graph as in Figure 14 (left). Then let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by repeating this procedure several times on the same graph, such that the clasps are “stacked” on top of each other, as indicated in Figure 14 (right). Note that the resulting graph can also be obtained by taking the vertex sum of $n$ copies of the original graph. In this sense, we can view these vertex sums as coming from clasp moves.
Fact 5.11.
Let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by taking the vertex sum of $n$ copies of the graph obtained from the clasp move in Figure 14 (left) on top of each other, as in Figure 14 (right). Then the bridge index of $G_{n}$ is less than or equal to $12n+$ 1.
@@
This family $G_{n}$ fits into a more general family of almost unknotted $\Theta_{m}$-graphs, constructed analogously, but we expect the bridge index to remain $12n+$ 1 regardless of $m$.
@@
We can construct a somewhat simpler family of almost unknotted $\Theta$-graphs by performing “half” of the clasp move in the initial graph, as in Figure 15 (left). Note that this graph fits into the family of Kinoshita-Wolcott graphs [kinoshita1958alexander, kinoshita1972elementary, wolcott1987knotting] (see [jang2016new, Fig. 1]), and could be seen as being obtained from this perspective. Let $G_{n}$ be the family obtained from taking $n$ vertex sums of this graph, as indicated in Figure 15 (right).
@@
In Figure 16 we give four consistent labelings of $G_{1}$ by the Alexander quandle of order four from Section 4.1. It is well-known that this quandle is homogeneous, but the readers can also verify from the table before Definition 4.2 that for any $x$ and $z$, there is a $y$ such that $x\triangleright y=z$. Observe that by permuting labels, the total number of colorings of $G_{1}$ is 16. By Proposition 4.4, the number of colorings of $G_{n}$ is $4^{n+1}$, and by a combination of the lower bound coming from Proposition 4.3 and the upper bound coming from our algorithm, we see that the bridge index is precisely $n+1$.
Corollary 5.13.
Let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by taking the vertex sum of $n$ copies of the graph obtained from the clasp move in Figure 15 (left) on top of each other, as in Figure 15 (right). Then the bridge index of $G_{n}$ equals $n+1$.
@@
Figure 16: Four consistent labelings of the graph in Figure 15 (left) by the Alexander quandle of order four. Thus by permuting the labels, we have a total of nine colorings.

### v2

@@ 4.2 Meridional rank
It is a straightforward exercise to see that any group produces a quandle structure given by $a\triangleright b=b^{-1}ab$. In fact, one motivation for quandle theory is to generalize the notion of conjugation in group theory. In the previous subsection, we counted the number of quandle colorings to get a bound on the number of seeds. When the quandle comes from a group, and the group is well-understood in terms of how to generate it, it suffices to find just one coloring (we do not have to count all possible colorings) to get a bound on the number of seeds. This is what we do in this subsection.
A presentation for the fundamental group of the complement of a spatial graph $G$ in $S^{3}$ can be computed via the Wirtinger algorithm in a similar way to the computations for knots and links (this can be proven using Van Kampen’s theorem). Each arc of a spatial graph diagram gives a generator, and each crossing gives a relation of the form $x_{i}x_{j}x^{-1}_{i}x^{-1}_{k}$. At a vertex, we have a new type of relation of the form $x_{i_{1}}^{\epsilon_{1}}x_{i_{2}}^{\epsilon_{2}}\cdots x_{i_{n}}^{\epsilon_{n}}$, where the generators are listed in order as we go clockwise around a vertex, and $\epsilon_{i}$ is 1 (resp. 1-1) if the arc is directed into the vertex (resp. directed out of the vertex). See Figure 16.
@@
The technique of bounding the number of handles of knotted objects by the meridional rank of the fundamental group has been used before in other settings such as classical links [BBK21] and surface-links [JP24]. We now adapt the method to the setting of spatial graphs.
[…]
@@
Recall that a presentation of a Coxeter group $C(\Delta)$ can be obtained from a weighted graph $\Delta$ as follows. Label the vertices of $\Delta$ as $x_{1},x_{2},\ldots,x_{n}$, which will be the generators. If there is an edge with weight $w$ connecting $x_{i}$ to $x_{j}$, then we have a relation $(x_{i}x_{j})^{w}$ in the presentation. Additionally, the relations also include $x_{i}^{2},$ for $i=1,2,\ldots,n.$
@@
Lemma 4.11 ([FT10, Lem. 2.1]).
If all edge weights of $\Delta$ are at least $2$, then the reflection rank equals the minimal number of reflections needed to generate $C(\Delta).$
@@
A labeling of arcs in a spatial graph diagram $D$ by a Coxeter group is consistent if at each crossing where the overstrand is labeled with an element $g$ and the two understrands are labeled $h$ and $k$, we have that $ghg=k$. We say that the label generates the group if every element of $C(\Delta)$ can be written as a product of the elements that appear as labels along with their inverses.
@@
If a diagram $D$ representing a spatial graph $G$ can be labeled by $n$ reflections from a Coxeter group $C(\Delta)$ with reflection rank $n$, then $\widehat{\beta}(G)\geq n.$
@@
In the same spirit as Proposition 4.6, the labels on seeds determine all the other labels. The difference is that we use a labeling by reflections in Coxeter groups in this setting. Thus, these labels generate $C(\Delta)$. We then appeal to Lemma 4.11 to conclude that we need at least $n$ reflections to generate $C(\Delta)$. ∎
@@ 5 Computations
In this section, we use our techniques to estimate or provide exact values for the bridge indices of many different examples of knotted spatial graphs, many of which are almost unknotted. Additionally, in Section 5.1 we discuss the particulars of our code and in Section 5.5 we prove Theorem 1.2. Refer to Section 2.1 for a review of the spatial graph terminology we use throughout this section. We will use the term “coloring” for both Wirtinger coloring and quandle coloring, but the type of coloring should be clear from context. When calculating lower bounds, we will sometimes refer to both the (weighted) bridge index and the unweighted bridge index (recall Definition 4.4).
We break into various subsections according to the type of construction, but start first with a relatively simple example.
@@
As a warmup, Figure 17 gives an almost unknotted handcuff graph (left) and an almost unknotted $\Theta$-graph (right). Both of these well-known examples – for instance the $\Theta$-graph is the Kinoshita graph [Kin58, Kin72] – were constructed by adding an edge to an unlink, and both have bridge index $\tfrac{5}{2}$. These diagrams are already in bridge position demonstrating this index, but this can also be verified by finding $\tfrac{5}{2}$-Wirtinger colorings of the diagrams.
Since there is a nontrivial coloring by the Alexander quandle of order four from Section 4.1, the unweighted bridge index, and thus the bridge index, is bounded below by $2$. By analyzing the various possibilities for how the vertices are placed in the bridge splitting, à la the arguments found in Example 5.8 or Example 5.14, one can – without too much difficulty – raise the lower bound of the bridge index to $\tfrac{5}{2}$, as desired.
Note that the number of crossings in each column of twists can be systematically increased and we still retain the consistency of the coloring. By systematic, we mean that if there are three crossings originally, we can keep adding a multiple of three crossings.
@@
Figure 17: A quandle colored almost unknotted handcuff graph (left) and a quandle colored almost unknotted $\Theta$-graph (right), both with bridge index $\tfrac{5}{2}$.
@@ 5.1 Python code
We have implemented our algorithm in Python to compute upper bounds for the Wirtinger numbers of arbitrary (knotted) $\Theta_{n}$-graphs ($n\leq 26)$. Our algorithm generalizes coloring algorithms used to compute bounds for the Wirtinger numbers for knots and links. We refer readers to [Bla+20a, LPV24] for details on the knot and link cases. Our Python code can be found at https://github.com/hanhv/graph-wirt. While the full detail can be found in the URL, we also give a summary here.
@@
A Gauss code for an $m$-component link is a list of $m$ lists of positive and negative integers such that for all $k\neq 0,|k|\leq m$, both $k$ and k-k appear once each among the lists.
@@
Figure 18: Example of a Gauss code. This diagram is $\tfrac{7}{2}$-Wirtinger colorable using the seeds [[a1, b1, c1], [-10, -13], [-9, 8, -3]], which are colored blue, green, and red respectively.
@@
First we extract two-component link diagrams from the link table provided by SnapPy [Cul+24]. Each of these links takes the form of a Gauss code made up of two lists of numbers. Selecting a two-component link, by “singularizing” we turn two crossings between components into 4-valent vertices. Here, a pair of crossings $(x,y)$ is considered singularizable if both $x$ and $y$ appear in one component, while x-x and y-y appear in the other component. We only choose singularizable pairs that are of distance at least four apart (this is optional), ensuring that each edge of the graph contains at least four crossings. By singularizing a pair of crossings $i$ and $j$, we split each component of the link, viewed as a cyclic list, into two parts with endpoints at $i$ and $j$. This has the effect of turning the two-component link into a $\Theta_{4}$-graph.
@@
In the context of our code, it is convenient to truncate our spatial graph diagrams, meaning we imagine removing small neighborhoods of vertices (of degree larger than two). We refer to these neighborhoods as pods. After removing pods, we have two types of arcs: arcs which connect the endpoints of pods to undercrossings, and arcs which connect two undercrossings.
The program begins by identifying all strands in the Gauss code which can be classified as pods or one of the two aforementioned types of arcs. It first sets $k=1$ and generates a list of seeds, [^4] Note that elsewhere in the paper we refer to each initially colored strand as a seed, but in the context of our code, we refer to the entire collection of initial strands as a seed. which are combinations of $k$ strands, ensuring that each strand must be either a pod or an arc which connects two undercrossings (more precisely, arcs connecting the endpoints of pods to undercrossings are not allowed to be in a seed). For each seed, it colors the strands in the seed and creates a truncated version of the graph by removing any pods that are not part of the selected seed.
The program then attempts to maximally extend this truncated graph through two types of coloring moves: one that extends from the endpoint of a pod and another that extends from an undercrossing if its corresponding overstrand is colored. If the entire truncated graph can be colored, the seed is added to the list of colorable seeds. Half of the sum of weights of strands in the seeds gives an upper bound for the Wirtinger number and an upper bound for $k$ (the maximum number of strands in a colorable seed). The program keeps running until $k$ reaches its upper bound and returns the minimum Wirtinger number and the corresponding seeds.
@@
We extracted and computed the Wirtinger numbers for over 10,000 $\Theta_{4}$-graph diagrams.
@@ 5.2 Hara’s graphs and modifications
The family of almost unknotted $\Theta_{4}$-graphs given by Hara [Har91, Fig. 2] all have unweighted bridge index $2$, and (weighted) bridge index at most $3$, which can be quickly verified by finding Wirtinger colorings of the diagrams (using the same seeds for both versions of bridge index). This is perhaps surprising, as these particular diagrams have many more local maxima and minima.
@@
For the sake of simplicity, we depict both families as coming from $\Theta$-graphs rather than $\Theta_{4}$-graphs, but both could be further extended by adding more edges in analogous ways. This would likely not change the bridge index. Note that with this change, we can (perhaps more appropriately) think of these families as modifications of a Kinoshita-Wolcott graph [Kin58, Kin72, Wol87] (see [Jan+16, Fig. 1]).
@@
Let $G_{n}$ be the $\Theta$-graph obtained from Hara’s graphs by replacing the braids on two strands with braids on $n$ strands. Then the unweighted bridge index $\widehat{\beta}(G_{n})$ is less than or equal to $n$, and the bridge index $\beta(G_{n})$ is less than or equal to $\tfrac{2n+1}{2}$.
@@
Let $G_{n}$ be the $\Theta_{3n}$-graph obtained from Hara’s graphs by taking $n$ parallel copies of each edge. Then the unweighted bridge index $\widehat{\beta}(G_{n})$ is less than or equal to $n+1$, and the bridge index $\beta(G_{n})$ is less than or equal to $\tfrac{5n}{2}$.
@@ 5.3 Families from vertex sums
Take $n$ copies of the almost unknotted Suzuki $\Theta_{4}$-graph [Suz84], as shown in Figure 20. Consider the quandle $(\{0,1,2\},\triangleright)$, where $x\triangleright y=2y-x\pmod{3}$. It can be shown that this quandle is homogeneous (see [FT24], for example). In Figure 20 we give three consistent colorings of the graph by this quandle; then by permuting the colors, we have a total of nine possible colorings.
Taking the vertex sum of $n$ copies of the graph, we obtain a diagram with bridge index $n+1$. Using Proposition 4.6, we get that the number of colorings of the resulting vertex sum is $3^{n+1}$. Then by Proposition 4.5, the minimum number of seeds is precisely $n+1$, where in this case each trivial tangle contains a vertex. Thus we have shown the following.
@@
Let $G_{n}$ be the $\Theta_{4}$-graph obtained by taking the vertex sum of $n$ copies of Suzuki’s graph. Then the unweighted bridge index $\widehat{\beta}(G_{n})$ is $n+1$.
With generalized 3-coloring (recall Definition 4.7 and Proposition 4.8), we can give a lower bound for the (weighted) bridge index $\beta$, not just the unweighted version $\widehat{\beta}$. The other type of bridge splitting that we have to consider is the case where two vertices of the $\Theta_{4}$-curve are on the same side. That is, one trivial tangle is a collection of intervals with no vertices.
[…]
2pt
[…]
Figure 20: For a Suzuki $\Theta_{4}$-graph, we see that the number of colorings where the arcs next to a particular vertex are labeled by a fixed element of the dihedral quandle of order three is three. (Here we use colors to represent the labelings by the quandle.) Thus by permuting the colors, we have a total of nine colorings. The labels in the leftmost diagram are used in the proof of Proposition 5.10.
Proposition 5.10.
The number of generalized $3$-colorings of the Suzuki $\Theta_{4}$ graph is $81=3^{4}$.
Proof.
We get the following chains of equations, referencing the labels in the leftmost diagram of Figure 20.
$y-z+t-w$ $=0\pmod{3}$ $a-b+c-d$ $=0\pmod{3}$
$2y-x$ $=a\pmod{3}$ $2a-y$ $=f\pmod{3}$
$2z-f$ $=b\pmod{3}$ $2b-z$ $=g\pmod{3}$
$2t-g$ $=c\pmod{3}$ $2c-t$ $=h\pmod{3}$
$2w-h$ $=d\pmod{3}$ $2d-w$ $=x\pmod{3}$
These equations can be transformed to the following.
$a$ $=2y-x\pmod{3}$ $b$ $=2z-x\pmod{3}$
$c$ $=2t-x\pmod{3}$ $d$ $=2w-x\pmod{3}$
$f$ $=g=h=x$
In other words, any element of $\{0,1,2\}$ is a valid choice for $x,y,z,$ and $t$, giving $3^{4}$ total number of solutions. ∎
This means that any bridge position in which one trivial tangle is made only of intervals requires at least four arcs in the trivial tangle.
Corollary 5.11.
Let $G_{n}$ be the $\Theta_{4}$-graph obtained by taking the vertex sum of $n$ copies of Suzuki’s graph. Then the bridge index $\beta(G_{n})$ is $n+2$.
Proof.
By Corollary 5.9, we have that the number of connected components in each tangle cannot be less than $n+1,$ but we have not ruled out the possibility that $\beta(G)=n+1.$ For that to happen, we must have that one of the trivial tangles is made only of intervals. However, Proposition 5.10 says that $G_{1}$ needs at least four arcs in a trivial tangle. Furthermore, when a vertex sum is performed, we can ensure that the colors at the strands where the sum happens have the same matching colors. For any fixed color at such a strand of one summand, there are three possible colored diagrams. Therefore, the number of generalized 3-coloring of $G_{2}=G_{1}\#_{4}G_{1}$ is $3^{4}\cdot 3=3^{5}$. Continuing with this logic, we conclude that the number of generalized 3-colorings of $G_{n}$ is $3^{n+3},$ implying that we need at least $n+3$ arcs of degree 2. So the (weighted) bridge index $\beta$ using only arcs in one trivial tangle is at least $n+3,$ which is strictly higher than bridge positions where vertices of the graphs are on different sides of the bridge splitting. ∎
@@
The clasp move [SW90] (see [Jan+16, Fig. 3]), is a well-known move which preserves almost unknottedness of $\Theta$-graphs. We will use this to build a family of graphs as follows. First construct an almost unknotted $\Theta$-graph by performing the clasp move on a slightly isotoped trivial $\Theta$-graph as in Figure 21 (left). Then let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by repeating this procedure several times on the same graph, such that the clasps are “stacked” on top of each other, as indicated in Figure 21 (right). Note that the resulting graph can also be obtained by taking the vertex sum of $n$ copies of the original graph. In this sense, we can view these vertex sums as coming from clasp moves.
We omit the proof of the following corollary, as it follows using either a very similar argument as for Corollary 5.15, or using the fact that this example can be seen as being obtained from vertex sums of the following, plus results in [TT18] and [TT21] about additivity of bridge index under vertex sum.
Corollary 5.13.
Let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by taking the vertex sum of $n$ copies of the graph obtained from the clasp move in Figure 21 (left) on top of each other, as in Figure 21 (right). Then the unweighted bridge index $\widehat{\beta}(G_{n})$ equals $2n+1$, and the bridge index $\beta(G_{n})$ equals $\tfrac{4n+3}{2}$.
@@
This family $G_{n}$ fits into a more general family of almost unknotted $\Theta_{m}$-graphs, constructed analogously, but we expect the bridge index to not change as $m$ increases.
@@
We can construct a somewhat simpler family of almost unknotted $\Theta$-graphs by performing “half” of the clasp move in the initial graph, as in Figure 22 (left). Note that this graph fits into the family of Kinoshita-Wolcott graphs [Kin58, Kin72, Wol87] (see [Jan+16, Fig. 1]), and could be seen as being obtained from this perspective. Let $G_{n}$ be the family obtained from taking $n$ vertex sums of this graph, as indicated in Figure 22 (right). Similarly, the graph obtained in Figure 21 (left) could be seen as being obtained by taking the vertex sum of the Kinoshita graph and its mirror.
@@
In Figure 23 we give four consistent labelings of $G_{1}$ by the homogeneous Alexander quandle of order four from Section 4.1. Observe that by permuting labels, the total number of colorings of $G_{1}$ is 16. By Proposition 4.6, the number of colorings of $G_{n}$ is $4^{n+1}$, and by a combination of the lower bound coming from Proposition 4.5 and the upper bound coming from our algorithm, we see that the unweighted bridge index is precisely $n+1$.
@@
Figure 23: Four consistent labelings of the graph in Figure 22 (left) by the Alexander quandle of order four. Thus by permuting the labels, we have a total of sixteen colorings.
To determine the (weighted) bridge index, we consider possibilities for the structure of the trivial tangles that fit in a bridge splitting of a $\Theta$-graph with a prescribed number $k$ of unknotted tree components in the upper tangle: there are four cases (see Figure 24 for $k=4$). Note that the configurations on the right column of Figure 24 will never realize the bridge index, so we only need to consider the left column configurations. That is, the bridge index is either $\frac{2n+3}{2}$ or $\frac{2n+2}{2}$. However, the latter cannot happen because the lower trivial tangle is forced to have $n$ connected components, contradicting our lower bound of $n+1$ from the quandle coloring.
Figure 24: Different possibilities for trivial tangles in a bridge splitting of a $\Theta$-graph, where the number of components for the tangle above is four.
Corollary 5.15.
Let $G_{n}$ be the almost unknotted $\Theta$-graph obtained by taking the vertex sum of $n$ copies of the graph obtained from the clasp move in Figure 22 (left) on top of each other, as in Figure 22 (right). Then the unweighted bridge index $\widehat{\beta}(G_{n})$ equals $n+1$, and the bridge index $\beta(G_{n})$ equals $\tfrac{2n+3}{2}$.
Our procedure for bounding the bridge index from below in this example yields the same lower bound previously achieved in [TT21]. However, their bounds (in Theorems 7.4 and 7.5) include hypotheses such that their results can only be applied to the present example (and by association, Example 5.1 and Example 5.12) in this paper.